# COHESION: Composite Graph Convolutional Network with Dual-Stage Fusion for Multimodal Recommendation

### v1

@@ Abstract.
Recent works in multimodal recommendations, which leverage diverse modal information to address data sparsity and enhance recommendation accuracy, have garnered considerable interest. Two key processes in multimodal recommendations are modality fusion and representation learning. Previous approaches in modality fusion often employ simplistic attentive or pre-defined strategies at early or late stages, failing to effectively handle irrelevant information among modalities. In representation learning, prior research has constructed heterogeneous and homogeneous graph structures encapsulating user-item, user-user, and item-item relationships to better capture user interests and item profiles. Modality fusion and representation learning were considered as two independent processes in previous work. In this paper, we reveal that these two processes are complementary and can support each other. Specifically, powerful representation learning enhances modality fusion, while effective fusion improves representation quality. Stemming from these two processes, we introduce a COmposite grapH convolutional nEtwork with dual-stage fuSION for the multimodal recommendation, named COHESION. Specifically, it introduces a dual-stage fusion strategy to reduce the impact of irrelevant information, refining all modalities using behavior modality (ID embedding) in the early stage and fusing their representations at the late stage. It also proposes a composite graph convolutional network that utilizes user-item, user-user, and item-item graphs to extract heterogeneous and homogeneous latent relationships within users and items. Besides, it introduces a novel adaptive optimization to ensure balanced and reasonable representations across modalities. Extensive experiments on three widely used datasets demonstrate the significant superiority of COHESION over various competitive baselines.
@@
journalyear: 2025copyright: acmlicensedconference: The 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–17, 2025; Padua, Italybooktitle: The 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR 2025), July 13–17, 2025, Padua, Italycopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systems;
@@ 1. Introduction
The rapid development of the Internet has led to information explosion, making recommender systems an indispensable tool in human society. Traditional recommender systems rely on modeling user preferences through historical user-item interactions (Koren et al., 2021; Xu et al., 2024d; Chen et al., 2025; Xu et al., 2024a). However, the data sparsity problem always frustrates the accuracy of recommendations. As the types of information on social media become increasingly diverse, multimodal information has recently been used to alleviate the data sparsity problem (Lei et al., 2023; Mu et al., 2022; Xu et al., 2024c). A line of work (He and McAuley, 2016; Chen et al., 2017) integrates multimodal information as side information to enhance item representations. To better learn the representation, many works (Wang et al., 2021; Wei et al., 2019; Xu et al., 2024b) construct the user-item bipartite graph and utilize the graph convolution network (GCN) to enhance the representation learning in multimodal recommendations. Despite some previous works achieving notable recommendation performance via constructing the heterogeneous user-item graph, the relationships between user-user and item-item remain unexplored. Recently, DualGNN (Wang et al., 2021) constructs the user-user graph to explore the hidden preference pattern by similar users. LATTICE (Zhang et al., 2021) builds the item-item graph to capture semantically correlative signals. Both of these two homogeneous graphs improve the performance of recommendations. In this paper, we pose two comprehensive questions:
Would exploiting both user-user and item-item graphs
achieve even better performance? And how?
@@
To answer this question, we select the most standard GCN-based multimodal recommendation model MMGCN [^1] MMGCN is the earliest and simplest graph-based multimodal recommendation model, which only contains GCNs without extra components. (Wei et al., 2019) and carefully design its three variants [^2] All of the user-user graph, item-item graph, early fusion strategy, and late fusion strategy in this investigation are designed following a recent survey paper (Zhou et al., 2023c).: MMGCNu, MMGCNi, and MMGCNui, which incorporate only user-user, only item-item, and both user-user and item-item graphs respectively. To assess the impact of fusion strategies, we implement the early fusion and late fusion strategies for each MMGCN variant, namely MMGCNe and MMGCNl, respectively. Table 1 shows the recommendation performance of eight MMGCN variants on the Baby dataset. In Table 1, our findings include: 1) The late fusion strategy outperforms the early fusion strategy when integrating only user-user or item-item graphs. Conversely, the early fusion strategy is more effective than late fusion when both graph types are integrated, possibly because late fusion permits deeper modality-specific learning, but it is more susceptible to the adverse effects of irrelevant inter-modality information. 2) Among the early fusion strategies, MMGCNui shows performance superior to MMGCNu and MMGCNi, suggesting that combining both user and item graphs facilitates more effective representation learning. These observations suggest that suboptimal fusion strategies in prior studies may have constrained the representation learning capabilities of the otherwise well-designed models. This hypothesis is empirically supported by the results presented in Table 6.
@@
To liberate the representation learning ability of composite graph convolutional network, we introduce a COmposite grapH convolutional nEtwork with dual-stage fuSION for a multimodal recommendation, named COHESION. Specifically, we propose a novel dual-stage fusion strategy to mitigate the negative effects of irrelevant information among different modalities by refining all modalities using behavior modality in the early stage and fusing all modalities’ representations in the late stage. We believe that refining all modalities using behavior modality can enable each modality to map to a better semantic space. This is because the goal of multimodal recommendation is to predict user preferences instead of understanding multimodal content. Moreover, we propose the composite graph convolutional network to learn the representations for all modalities, including homogeneous and heterogeneous graphs. More specifically, we utilize a user-item graph to mine the relations between users and items and utilize user-user and item-item graphs to extract latent relations within users and items, respectively. Thanks to our dual-stage fusion strategy, our composite graph convolutional network liberates outstanding representation learning ability. Finally, we fuse all modality representations in the late stage and propose an adaptive optimization, which can make the representations learned from different modalities balanced. Extensive experiments on three widely used datasets demonstrate the superiority of COHESION over various competitive baselines. The contributions of our work are as follows:
  • We reveal the influences of fusion strategies to representation learning in multimodal recommendations.
  • We design a new composite graph convolutional network with a dual-stage fusion strategy, which liberates the representation learning ability.
  • We introduce a plug-and-play adaptive optimization, which enhances the overall learning effect for all modalities through adaptive weights.
  • Extensive experiments on three public datasets show the effectiveness of our COHESION.
@@ 2.2. Multimodal Fusion
With the rapid development of representation learning ability in recent multimodal recommendation models, predefined fusion strategies are more fragile to the negative impact of irrelevant information among modalities and are difficult to fuse different modalities effectively. To alleviate these problems, we propose a tailored dual-stage fusion strategy, which effectively liberates the representation learning ability for multimodal recommendation models.
@@ 3. Methodology
Figure 1. The overview structure of COHESION. It first proposes a novel early fusion strategy, which reduces the negative effect of irrelevant information among modalities by refining all modalities using behavior modality. Then it utilizes heterogeneous user-item GCN to learn the modality-specific representations. After that, a late fusion strategy is applied to fuse all the modalities. The representations of users and items are enhanced with homogeneous graphs. Finally, we propose an adaptive optimization to enhance modality-specific learning.
In this section, we present our proposed model COHESION [^3] Code is available at: https://github.com/Jinfeng-Xu/COHESION. in detail. The overall architecture is shown in Fig. 1.
@@ 3.1. Preliminary
Let a set of users $u\in\mathcal{U}$ and a set of items $i$ = $\{i_{m}|i_{i},i_{t},i_{v}\}\in\mathcal{I}$, where $m\in\mathcal{M}$ is the modality, $\mathcal{M}$ is the set of modalities, and $i_{i}$, $i_{t}$, and $i_{v}$ represent the behavior, textual, and visual modality features, respectively. $\mathcal{G}$ = $(\mathcal{V},\mathcal{E})$ be a given graph with a node set $\mathcal{V}$ and an edge set $\mathcal{E}$, where $|\mathcal{V}|$ = $|\mathcal{U}|$ + $|\mathcal{I}|$. The user-item interaction matrix is denoted as $\mathcal{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|}$. Specifically, each entry $\mathcal{R}_{u,i}$ indicates whether the user $u$ is connected to item $i$, with a value of 1 representing a connection and 0 otherwise. The graph structure of $\mathcal{G}$ can be denoted as the adjacency matrix $\mathcal{A}\in\mathbb{R}^{(|\mathcal{U}|+|\mathcal{I}|)\times(|\mathcal{U}|+|\mathcal{I}|)}$, formally:
@@ 3.2. Early Fusion
where $\dagger{E}_{i_{m}^{i}}$ denotes refined modality embedding for modality $m$, and $\epsilon$ is a positive infinitesimal to prevent zero vector. Through this tailored yet simple early fusion strategy, we enhance the representation learning capability of multimodal recommendation models by refining different modalities via the behavior modality. This early fusion strategy mitigates the gap between modalities, leading to more accurate and robust recommendations
@@ 3.5.2. Item-Item Graph
where $A^{(l)}_{i}$ is $l$-th layer representation of item $i$ learned from $\mathcal{G}^{i}$. $A^{(0)}_{i}$ is the item side $E_{f}(I)$ of $\dagger{E}_{f}$, which can be obtained by $\dagger{E}_{f}[|\mathcal{U}|:]$.
Then, we enhance the user and item representations by user-user and item-item graphs, respectively:
@@ 4.1.2. Baseline
To evaluate the effectiveness of our COHESION, we compare it with the following methods, which can be divided into two groups. The first group is the traditional recommendation models, including MF-BPR (Rendle et al., 2012), LightGCN (He et al., 2020), SimGCL (Yu et al., 2022), and LayerGCN (Zhou et al., 2023a). The second group is the multimodal recommendation models, including VBPR (He and McAuley, 2016), MMGCN (Wei et al., 2019), DualGNN (Wang et al., 2021), LATTICE (Zhang et al., 2021), FREEDOM (Zhou and Shen, 2023), SLMRec (Tao et al., 2022), BM3 (Zhou et al., 2023b), and LGMRec (Guo et al., 2024). We detailed all them as follows:
@@
  • MF-BPR (Rendle et al., 2012) utilizes the Bayesian Personalized Ranking (BPR) loss to optimize the classic collaborative filtering method, which learns user and item representations by matrix factorization method.
@@
  • SimGCL (Yu et al., 2022) proposes a graph contrastive learning method, which directly injects random noises into representation.
@@
  • LATTICE (Zhang et al., 2021) proposes an item-item graph to capture semantically correlative signals among items.
@@
  • LGMRec (Guo et al., 2024) captures and utilizes local embeddings with local topological information and global embeddings with hypergraphs.
@@ 4.4. Effectiveness of Mitigating Irrelevant Information among Modalities (RQ3)
Figure 3. Visualization of modality distribution through t-SNE. We randomly sample 500 items from Baby dataset. Red, green, and blue represent the ID, textual, and visual modalities.
@@ 4.5. Different Sparse Data Scenarios (RQ4)
Table 7. Comparison of our COHESION against state-of-the-art baselines on efficiency. Parameter denotes the learnable parameter for the model, and Time denotes the total duration of each iteration, including both the training and evaluation phases.

### v2

@@ Abstract.
Recent works in multimodal recommendations, which leverage diverse modal information to address data sparsity and enhance recommendation accuracy, have garnered considerable interest. Two key processes in multimodal recommendations are modality fusion and representation learning. Previous approaches in modality fusion often employ simplistic attentive or pre-defined strategies at early or late stages, failing to effectively handle irrelevant information among modalities. In representation learning, prior research has constructed heterogeneous and homogeneous graph structures encapsulating user-item, user-user, and item-item relationships to better capture user interests and item profiles. Modality fusion and representation learning were considered as two independent processes in previous work. This paper reveals that these two processes are complementary and can support each other. Specifically, powerful representation learning enhances modality fusion, while effective fusion improves representation quality. Stemming from these two processes, we introduce a COmposite grapH convolutional nEtwork with dual-stage fuSION for the multimodal recommendation, named COHESION. Specifically, it introduces a dual-stage fusion strategy to reduce the impact of irrelevant information, refining all modalities using behavior modality in the early stage and fusing their representations at the late stage. It also proposes a composite graph convolutional network that utilizes user-item, user-user, and item-item graphs to extract heterogeneous and homogeneous latent relationships within users and items. Besides, it introduces a novel adaptive optimization to ensure balanced and reasonable representations across modalities. Extensive experiments on three public datasets demonstrate the significant superiority of COHESION over various competitive baselines.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729927isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems;
@@ 1. Introduction
The rapid development of the Internet has led to information explosion, making recommender systems an indispensable tool in human society. Traditional recommender systems rely on modeling user preferences through historical user-item interactions (Koren et al., 2021; Xu et al., 2024c; Chen et al., 2025; Xu et al., 2024a). However, the data sparsity problem always frustrates the accuracy of recommendations. As the types of information on social media become increasingly diverse, multimodal information has recently been used to alleviate the data sparsity problem (Lei et al., 2023; Mu et al., 2022; Xu et al., 2024d). A line of work (He and McAuley, 2016; Chen et al., 2017) integrates multimodal information as side information to enhance item representations. To better learn the representation, many works (Wang et al., 2021; Wei et al., 2019; Xu et al., 2024b) construct the user-item bipartite graph and utilize the graph convolution network (GCN) to enhance the representation learning in multimodal recommendations. Despite some previous works achieving notable recommendation performance via constructing the heterogeneous user-item graph, the relationships between user-user and item-item remain unexplored. Recently, DualGNN (Wang et al., 2021) constructs the user-user graph to explore the hidden preference pattern by similar users. LATTICE (Zhang et al., 2021) builds the item-item graph to capture semantically correlated signals. Both of these two homogeneous graphs improve the performance of recommendations. In this paper, we pose two comprehensive questions: Would exploiting both user-user and item-item graphs achieve even better performance? And how? To answer this question, we select the most standard GCN-based multimodal recommendation model MMGCN [^1] MMGCN is the earliest and simplest graph-based multimodal recommendation model, which only contains GCNs without extra components. (Wei et al., 2019) and carefully design its three variants [^2] All of the user-user graph, item-item graph, early fusion strategy, and late fusion strategy in this investigation are designed following a recent survey paper (Zhou et al., 2023c).: MMGCNu, MMGCNi, and MMGCNui, which incorporate only user-user, only item-item, and both user-user and item-item graphs, respectively. To assess the impact of fusion strategies, we implement the early fusion and late fusion strategies for each MMGCN variant, namely MMGCNe and MMGCNl, respectively. Table 1 shows the recommendation performance of eight MMGCN variants on the Baby dataset. In Table 1, our findings include: 1) The late fusion strategy outperforms the early fusion strategy when integrating only user-user or item-item graphs. Conversely, the early fusion strategy is more effective than late fusion when both graph types are integrated, possibly because late fusion permits deeper modality-specific learning, but it is more susceptible to the adverse effects of irrelevant inter-modality information. 2) Among the early fusion strategies, MMGCNui shows performance superior to MMGCNu and MMGCNi, suggesting that combining both user and item graphs facilitates more effective representation learning. These observations suggest that suboptimal fusion strategies in prior studies may have constrained the representation learning capabilities of the otherwise well-designed models. This hypothesis is empirically supported by the results presented in Table 6.
@@
To liberate the representation learning ability of composite graph convolutional network, we introduce a COmposite grapH convolutional nEtwork with dual-stage fuSION for a multimodal recommendation, named COHESION. Specifically, we propose a novel dual-stage fusion strategy to mitigate the negative effects of irrelevant information among different modalities by refining all modalities using behavior modality in the early stage and fusing all modalities’ representations in the late stage. We believe that refining all modalities using behavior modality can enable each modality to map to a better semantic space. This is because the goal of multimodal recommendation is to predict user preferences instead of understanding multimodal content. Moreover, we propose the composite graph convolutional network to learn the representations for all modalities, including homogeneous and heterogeneous graphs. More specifically, we utilize a user-item graph to mine the relations between users and items and utilize user-user and item-item graphs to extract latent relations within users and items, respectively. Thanks to our dual-stage fusion strategy, our composite graph convolutional network liberates outstanding representation learning ability. Finally, we fuse all modality representations in the late stage and propose an adaptive optimization, which can make the representations learned from different modalities balanced. Extensive experiments on three widely used datasets demonstrate the superiority of COHESION over various competitive baselines.
@@ 2.2. Multimodal Fusion
With the rapid development of representation learning ability in recent multimodal recommendation models, predefined fusion strategies are more fragile to the negative impact of irrelevant information among modalities and are difficult to fuse different modalities effectively. To this end, we propose a tailored dual-stage fusion strategy, which effectively liberates the representation learning ability for multimodal recommendation models.
@@ 3. Methodology
Figure 1. The overview structure of COHESION.
In this section, we present our proposed COHESION [^1] The code will be released at: https://github.com/Jinfeng-Xu/COHESION after the SIGIR’25 conference. in detail. The overall architecture is shown in Fig. 1.
@@ 3.1. Preliminary
Let a set of users $u\in\mathcal{U}$ and a set of items $i$ = $\{i_{m}|i_{i},i_{t},i_{v}\}\in\mathcal{I}$, where $m\in\mathcal{M}$ is the modality, $\mathcal{M}$ is the set of modalities, and $i_{i}$, $i_{t}$, and $i_{v}$ represent the behavior, textual, and visual modality features, respectively. $\mathcal{G}$ = $(\mathcal{V},\mathcal{E})$ be a given graph with a node set $\mathcal{V}$ and an edge set $\mathcal{E}$, where $|\mathcal{V}|$ = $|\mathcal{U}|$ + $|\mathcal{I}|$. The user-item interaction matrix is denoted as $\mathcal{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|}$. Specifically, each entry $\mathcal{R}_{u,i}$ indicates whether the user $u$ is connected to item $i$, with a value of 1 representing a connection and 0 otherwise. The graph structure of $\mathcal{G}$ can be denoted as the adjacency matrix $\mathcal{A}\in\mathbb{R}^{(|\mathcal{U}|+|\mathcal{I}|)\times(|\mathcal{U}|+|\mathcal{I}|)}$:
@@ 3.2. Early Fusion
where $\dagger{E}_{i_{m}^{i}}$ denotes refined modality embedding for modality $m$, and $\epsilon$ is a positive infinitesimal to prevent zero vector. Through this tailored yet simple early fusion strategy, we enhance the representation learning capability of multimodal recommendation models by refining different modalities via the behavior modality. This early fusion strategy mitigates the gap between modalities, leading to more accurate and robust recommendations.
@@ 3.5.2. Item-Item Graph
where $A^{(l)}_{i}$ is $l$-th layer representation of item $i$ learned from $\mathcal{G}^{i}$. $A^{(0)}_{i}$ is the item side $E_{f}(I)$ of $\dagger{E}_{f}$, which can be obtained by $\dagger{E}_{f}[|\mathcal{U}|:]$. Then, we enhance the user and item representations by user-user and item-item graphs, respectively:
@@ 4.1.2. Baseline
To evaluate the effectiveness of our COHESION, we compare it with the following methods, which can be divided into two groups. We detailed all of them as follows:
@@
  • MF-BPR (Rendle et al., 2012) utilizes the BPR loss to optimize the classic collaborative filtering method, which learns user and item representations by the matrix factorization method.
@@
  • SimGCL (Yu et al., 2022) proposes a graph contrastive learning method, which directly injects random noises into the representation.
@@
  • LATTICE (Zhang et al., 2021) proposes an item-item graph to capture semantically correlated signals among items.
@@
  • LGMRec (Guo et al., 2024) utilizes local embeddings with local topological information and global embeddings with hypergraphs.
@@ 4.4. Effectiveness of Mitigating Irrelevant Information among Modalities (RQ3)
Figure 3. Visualization of distribution through t-SNE. We randomly sample 500 items from Baby dataset. Red, green, and blue represent the ID, textual, and visual modalities.
@@ 4.5. Different Sparse Data Scenarios (RQ4)
Table 7. Comparison of COHESION against competitive baselines on efficiency. Parameter denotes the learnable parameter for the model, and Time denotes the total duration of each iteration, including both the training and evaluation phases.